OpenStudy (anonymous):

-3 + 8x - 5 = -8 PLEASE :) help me solve for x

hartnn (hartnn):

what is -3-5 ?

OpenStudy (anonymous):

negative three plus eight x minus five equals negative eight

OpenStudy (anonymous):

since you want to find the value of x, transfer the non x terms to ur right side, so you'll get 8x=-8+5+5 now solve the ones on your right 8x=2 soo....now you can find x... x=2/8 it can be simplified somemore.. x=1/4. there you go

hartnn (hartnn):

lol! its not -5 its -3 8x=-8+5+3 8x=0 x=??

OpenStudy (anonymous):

woops, careless mistake. must avoid during exammm hahaha

OpenStudy (anonymous):

x=0